OpenStudy (pulsified333):

A box contains 4 red and 5 green marbles. If 5 marbles are drawn without replacement, what is the expected number of red marbles?

OpenStudy (anonymous):

multiply

OpenStudy (pulsified333):

@satellite73 Multiply what?

Nnesha (nnesha):

total number of marbles = 9 4 are red so first draw = 4/5 2nd draw 3/8 ( we already picked one red marble that gives us 4-1 =3 subtract one from total number of marbles 9-1= 8) 3rd draw = 2/7 ^ 4th = 1/6 5th = 0/5 now multiply all fraction i hope that's correct :D

Nnesha (nnesha):

oh wait no :D

OpenStudy (pulsified333):

on the first draw how is it 4/5?

Nnesha (nnesha):

typo 4/9**

Nnesha (nnesha):

hmm that looks incorrect hmm

OpenStudy (pulsified333):

@Nnesha so will it be right after the 4/9 correction?

Nnesha (nnesha):

hmm i'm not sure about 0/5 that will give us 0 as a final answer hmm

OpenStudy (pulsified333):

i got 24/3024 as the answer but that doesn't seem right because we aren't asking for probability

OpenStudy (anonymous):

ok i lied, this is more complicated that in looks

OpenStudy (anonymous):

the probability you get 0 red is 0 (one has to be red)

OpenStudy (anonymous):

the probabiliy you get exactly one red is \[\frac{\binom{4}{1}\binom{5}{4}}{\binom{9}{5}}\]

OpenStudy (pulsified333):

ok

OpenStudy (pulsified333):

then you do that for 2 and 3 and 4

OpenStudy (anonymous):

that is not really so hard to compute as it looks \[\frac{4\times 5}{126}\]

OpenStudy (anonymous):

yeah then repeat for 2, 3, 4 then multiply, then add

OpenStudy (anonymous):

1 times the probability you get 1, plus 2 times the probability you get 2 etc

OpenStudy (anonymous):

there may be a simpler way to do it but i don't know it i would cheat

OpenStudy (pulsified333):

@satellite73 for the denominator would it be going down by one every time because it says without replacement?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that is taken in to consideration when you use the formula

OpenStudy (pulsified333):

oh okay

OpenStudy (pulsified333):

why am i multiplying 2 times the probability of getting 2?

OpenStudy (anonymous):

the pobabilty you get for example 3 red and 2 white is \[\frac{\binom{4}{3}\binom{5}{2}}{\binom{9}{5}}\]

OpenStudy (anonymous):

because that is how you get an expected value the number you get times the probability you get it

OpenStudy (pulsified333):

oh

OpenStudy (anonymous):

\[P(x=1)+2 P(x=2)+3P(x=3)+4P(x=4)\]

OpenStudy (anonymous):

in this case

OpenStudy (anonymous):

more if there are more possibilities

OpenStudy (pulsified333):

I get a huge number when I do that

OpenStudy (anonymous):

i got this, but i wouldn't bet more that $7 on it not at all http://www.wolframalpha.com/input/?i=20%2F126%2B2%284+choose+2%29%285+choose+3%29%2F126%2B3%284+choose+3%29%285+choose+2%29%2F126%2B+4*5%2F126

OpenStudy (anonymous):

it is not huge

OpenStudy (anonymous):

we can do them one at a time if you like

OpenStudy (pulsified333):

that was right :D

OpenStudy (anonymous):

imagine!